The following graph represents the situation of Sindbads caps, a firm selling caps in the perfectly competitive caps industr1) How much output should Sindbad produce to maximize his profit, if the market price is equal
to $11? (2 marks)
2) How much profit (loss) will he earn? (2 marks)
3) Indicate the profit (loss) area on the graph. (2 marks)
4) Find the fixed cost paid by the firm. (2 marks)
5) Suppose Sindbad decides to shut down. What would his loss be? (2 marks)
Answer #1

1) How much output should Sindbad produce to maximize his profit, if the market price is equal
to $11?

Ans: 100 quantity

explanation: firm maximizes its profit where MR equals to MC. here at quantity 100, MR=MC.

2) How much profit (loss) will he earn?

Ans: $200.

Explanation:

profit=TR-TC

=1100-900

=200.

TR=P*Q

=11*100

=1100.

TC=ATC*Q

=9*100

=900.

3) Indicate the profit (loss) area on the graph.

MC psont Aic AVMR ENET - 50 So 100 Quantity

4) Find the fixed cost paid by the firm.

Ans: 300.

Explanation:

FC=TC-TVC

=900-600

=300.

TC=ATC*Q
=9*100

=900.

TVC=AVC*Q

=6*100

=600

5) Suppose Sindbad decides to shut down. What would his loss be?

Ans: 200

explanation: if he will shutdown loss will be total profit earn by firm, if it produce.